# Finte-size induced random switching of chimeras in a deterministic two-population Kuramoto-Sakaguchi model

### v1

@@ Abstract
The two-population Kuramoto-Sakaguchi model for interacting populations of phase oscillators exhibits chimera states whereby one population is synchronised and the other is desynchronised. Which of the two populations is synchronised depends on the initial conditions. We show that this deterministic model exhibits random switches of their chimera states, alternating between which of the two populations is synchronised and which is not. We show that these random switches are induced by the finite size of the network. We provide numerical evidence that the switches are governed by a Poisson process and that the time between switches grows exponentially with the system size, rendering switches unobservable for all practical purposes in sufficiently large networks. We develop a reduced stochastic model for the synchronised population, based on a central limit theorem controlling the collective effect of the desynchronised population on the synchronised one, and show that this stochastic model well reproduces the statistical behaviour of the full deterministic model. We further determine critical fluctuation sizes capable of inducing switches and provide estimates for the mean switching times from an associated Kramers problem.
@@ I Introduction
Coupled oscillators describe systems ranging from neurons in the brain Crook et al. (1997); Sheeba, Stefanovska, and McClintock (2008); Bhowmik and Shanahan (2012) and synchronous firefly flashing Mirollo and Strogatz (1990), to power grids Filatrella, Nielsen, and Pedersen (2008); Nishikawa and Motter (2015) and Josephson junction arrays Watanabe and Strogatz (1994); Wiesenfeld, Colet, and Strogatz (1998). The celebrated Kuramoto model and its numerous variants Kuramoto (1984); Sakaguchi and Kuramoto (1986); Strogatz (2000); Pikovsky, Rosenblum, and Kurths (2001); Acebrón et al. (2005); Osipov, Kurths, and Zhou (2007); Arenas et al. (2008); Dörfler and Bullo (2014); Rodrigues et al. (2016) constitutes a paradigmatic model to describe the temporal evolution of such coupled oscillators and its often rich dynamical behaviour. Much progress has been made by employing mean-field theory, valid in the thermodynamic limit of infinitely many oscillators Sakaguchi and Kuramoto (1986); Ott and Antonsen (2008); Delgadino, Gvalani, and Pavliotis (2021). However, recently a number of finite-size effects have attracted attention such as finite-size induced stochastic drifts of oscillators which disappear in the thermodynamic limit Giacomin and Poquet (2015); Luçon (2015); Gottwald (2015, 2017). Such finite-size effects are not amenable to the classical mean-field theory and require new analytical approaches Hildebrand, Buice, and Chow (2007); Buice and Chow (2007); Hong, Ha, and Park (2007); Tang (2011); Giacomin and Poquet (2015); Luçon (2015); Gottwald (2015, 2017).
@@
We present a dynamical phenomenon which is due entirely to the finite size of the number of oscillators. We describe the emergence of a stochastic switching of chimera states in a deterministic system of a two-population Kuramoto-Sakaguchi (KS) model Okuda and Kuramoto (1991); Montbrió, Kurths, and Blasius (2004). Chimeras denote a symmetry breaking state in which a group of oscillators is synchronised while the rest of the oscillators are desynchronised and where the partition into a synchronised and desynchronised group depends on the initial condition Kuramoto and Battogtokh (2002); Omel’chenko, Maistrenko, and Tass (2008); Omel’chenko, Wolfrum, and Maistrenko (2010); Omel’chenko (2018). The existence of chimera states in a two-population KS model is well known and has been studied in either small networks Panaggio et al. (2016); Burylko, Martens, and Bick (2022) or in the thermodynamic limit Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016). Chimeras can undergo complex dynamical behaviour. In switching (or alternating) chimeras the synchronised and desynchronised regions alternate in time. To date such switching has been attributed to either the existence of metastable states Ma, Wang, and Liu (2010); Laing (2012); Buscarino et al. (2015); Semenova et al. (2016); Zhang et al. (2020) or toheteroclinic cycles Bick (2018, 2019); Bick and Lohse (2019); Ebrahimzadeh et al. (2020); Goldschmidt, Pikovsky, and Politi (2019); Brezetsky et al. (2021); Lee and Krischer (2023). We report here on the effective stochastic behaviour of switching between a synchronised population and a desynchronised population in a deterministic two-population KS model, where the switching is mediated entirely by finite-size effects and disappears in sufficiently large networks. Numerical simulations will show that the mean switching time increases exponentially with the number of oscillators $N$ and that the variance of the order parameter scales as $1/N$, suggesting an underlying central limit theorem which captures fluctuations around the well-known deterministic thermodynamic limit. We will describe the fluctuations by means of the recently proposed stochastic model reduction for deterministic KS models Yue and Gottwald (2024). We adopt the stochastic modelling framework to parametrise the interaction term that couples the phases of the desynchronised population to those of the synchronised one by a Gaussian process. Gaussian processes are entirely determined by their mean and their covariance function. The mean, represented by the thermodynamic limit of infinitely many oscillators, can be estimated by classical mean field theory. The covariance function will be estimated by first approximating the Gaussian process by a complex Ornstein-Uhlenbeck process, and then numerically determining the parameters of this complex Ornstein-Uhlenbeck process to fit the observed covariance function of the finite-size fluctuations. This modelling strategy captures the statistical features of the full two-population KS model very well, including the distribution of the order parameter. We will further determine the critical fluctuation size which is sufficient to induce chimera switching, and set up a classical Kramers problem to approximate the mean switching time with reasonable accuracy Kramers (1940). Our modelling strategy requires the system size to be sufficiently large to allow for the underlying central limit theorem to be valid and sufficiently small to prevent the law of large numbers described by the thermodynamic limit and classical mean-field theory.
@@ II Kuramoto-Sakaguchi model for two interacting populations
To account for different populations of oscillators and their mutual interaction extensions of the KS model were proposed Okuda and Kuramoto (1991); Montbrió, Kurths, and Blasius (2004). The intricate balance between the interactions of oscillators belonging to different populations gives rise to rich dynamical behaviour. In particular, these systems support chimera states in which one population is synchronised while the other is desynchronised Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016); Panaggio et al. (2016); Burylko, Martens, and Bick (2022). Which of the two populations is synchronised and which is not depends on the initial condition. We consider here the dynamics of equally sized populations $P_{1}$ and $P_{2}$ of size $N_{1}=N_{2}=N$, and where each population is governed by the same dynamics in the sense that their native frequencies are drawn from the same distribution and their respective intra- and interpopulation interactions are of equal strength. The dynamics of the phase oscillators is given by
@@
$\frac{d}{dt}\theta_{j}^{(2)}$ $=\omega_{j}^{(2)}+\frac{K}{N}\sum_{j\in P_{2}}\sin(\theta^{(2)}_{j}-\theta^{(2)}_{i}-\lambda)$
@@
Here $\theta_{i}^{(\alpha)}$ denotes the phase of oscillator $i=1,\ldots,N$ in population $\alpha\in\{1,2\}$. The intrapopulation and interpopulation coupling strengths are denoted by $K$ and $\kappa$, respectively. We assume attractive forces and a stronger interaction between oscillators belonging to the same population with $K>\kappa>0$. The parameter $\lambda\in[0,\frac{\pi}{2}]$ accounts for a phase lag. Each oscillator is equipped with a native frequency drawn from a distribution $g(\omega)$. Note that through a change of coordinates into a rotating reference frame, $\theta_{i}^{(\alpha)}(t)\to\theta^{(\alpha)}_{i}(t)-\bar{\omega}\,t$ with mean native frequency $\bar{\omega}=\frac{1}{N}\sum\omega_{i}$, we may assume without loss of generality that the mean native frequency $\bar{\omega}$ is zero. Here we consider normally distributed natural frequencies $g(\omega)\sim\mathcal{N}(0,\sigma_{\omega}^{2})$ for both populations, and in all computations we set $\sigma_{\omega}^{2}=1$. In the following we shall use greek sub- or superscripts to label populations and roman subscripts to label indices of oscillators.
The two-population KS model (1)–(2) undergoes a transition from desynchronised behaviour to synchronisation as the coupling strengths $K$ and $\kappa$ are gradually increased. At lower coupling strengths the oscillators exhibit desynchronised dynamics and a lack of coherent phase alignment. With increasing coupling strengths, the oscillators align their phases and oscillate with a common rotation frequency $\Omega$. Notably, this synchronisation typically involves both populations achieving coherent phase locking. However, depending on the particular choice of the intrapopulation coupling strength $K$ and the interpopulation coupling strength $\kappa$, the system may exhibit symmetry breaking between the two populations and chimera states appear wherein one population achieves synchronisation while the other remains in a (relatively) desynchronised state Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016); Burylko, Martens, and Bick (2022).
The collective behaviour of each of the populations is captured by complex mean-field variables $z_{\alpha}$, or equivalently by the real mean-field variables $r_{\alpha}$ and $\psi_{\alpha}$ for $\alpha=1,2$, with
$z_{\alpha}\coloneq r_{\alpha}(t)e^{i\psi_{\alpha}(t)}=\frac{1}{N}\sum_{j\in P_{\alpha}}^{N}e^{i\theta^{(\alpha)}_{j}(t)}.$ (3)
@@
${\bar{r}}_{\alpha}=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}r_{\alpha}(t)dt,$
quantifies the degree of synchronisation within a population $P_{\alpha}$; complete phase synchronisation with $\theta^{(\alpha)}_{1}=\theta^{(\alpha)}_{2}=\cdots=\theta^{(\alpha)}_{N}$ implies $\bar{r}_{\alpha}=1$, whereas a fully desynchronised population with spread out phases implies $\bar{r}_{\alpha}\sim\mathcal{O}(1/\sqrt{N})$.
@@ II.1 Mean-field theory
Sakaguchi and Kuramoto Sakaguchi and Kuramoto (1986) developed a mean-field theory for the case of a single population with uniform coupling for the order parameter $r$ and the mean frequency $\Omega$ of the synchronised cluster. Mean-field theory has been extended for the two population KS model (1)–(2) Abrams et al. (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020). In the two-population KS model, we have mean-field variables for each population, i.e. $z_{\alpha}$, and the common mean frequency $\Omega$. The two-population KS model (1)–(2) can be written in terms of the complex mean field variables (3) as
$\frac{d}{dt}\theta^{(1)}_{i}$ $=\omega_{i}^{(1)}-\Omega+\textrm{Im}\left[(Kz_{1}+\kappa z_{2})e^{-i\lambda}e^{-i\theta^{(1)}_{i}}\right$] (4)
@@
In view of describing chimera states in which one of the two populations is synchronised whereas the other one is not, we shall without loss of generality assume that population $P_{1}$ is (partially) synchronised and $P_{2}$ is desynchronised, and we move into the frame of reference of the synchronised cluster with mean phase $\psi_{1}$ and mean frequency $\Omega$. This implies $z_{1}e^{-i\psi_{1}}=r_{1}$ and motivates to introduce the phase-shifted order parameter of the desynchronised population
$Z\coloneq z_{2}=r_{2}e^{i\Delta\psi},$ (6)
@@
$\frac{d}{dt}\theta^{(\alpha)}_{j}$ $=\omega_{j}^{(\alpha)}-\Omega+\text{Im}(v_{\alpha}e^{-i\theta_{j}})$ (7)
@@
In the thermodynamic limit $\Omega$, $r_{1}$ and $Z$ are constant. For constant mean-field parameters stationary solutions of (7)–(8) exist, provided the coupling strengths are sufficiently large ensuring
@@
$\theta^{(\alpha)}_{j}$ $=\arcsin(\frac{\omega^{(\alpha)}_{j}-\Omega}{|v_{\alpha}|})+\arg(v_{\alpha}).$ (10)
@@
In the thermodynamic limit $N\to\infty$, the phases for each population can be described by a probability density function $\rho^{(\alpha)}(\theta,t;\omega)$ for $\alpha=1,2$, satisfying the continuity equation
$\frac{\partial\rho^{(\alpha)}}{\partial t}+\frac{\partial}{\partial\theta}(\nu^{(\alpha)}\rho^{(\alpha)})=0,$ (11)
where the vectorfields $\nu^{(\alpha)}$ are given by the right-hand-side of (7) with
$\nu^{(\alpha)}$ $=\omega^{(\alpha)}-\Omega+\text{Im}(v_{\alpha}e^{-i\theta^{(\alpha)}}),$ (12)
Stationary solutions satisfying $\partial_{t}\rho^{(\alpha)}=0$ depend on the equation parameters. Entrained oscillators with frequencies satisfying (9) are distributed according to the stationary density
@@
$\rho^{(\alpha)}(\theta;\omega)$ $=C^{(\alpha)}\frac{1}{\nu^{(\alpha)}},$ (14)
with normalisation constant $C^{(\alpha)}$, such that the density is high in regions where the phases spend much of their time (cf.(7)–(8)).
The order parameters $r_{1}$ and $Z=r_{2}e^{i\Delta\psi}$ are given in the thermodynamic limit as
$r_{1}$ $=\int_{-\infty}^{\infty}\int_{0}^{2\pi}e^{i\theta}\rho^{(1)}(\theta,t)g(\omega)d\theta d\omega$ (15)
$Z$ $=\int_{-\infty}^{\infty}\int_{0}^{2\pi}e^{i\theta}\rho^{(2)}(\theta,t)g(\omega)d\theta d\omega.$ (16)
The integration over the frequencies has to be split into the entrained and non-entrained ranges (9), each with their respective stationary density (13) or (14). After some algebraic manipulation, we obtain
$r_{1}$ $=i\int_{-\infty}^{\infty}\frac{v_{1}}{|v_{1}|}\left[\frac{\omega-\Omega}{|v_{1}|}\right$.
$\left.\qquad-\sqrt{\frac{\omega-\Omega}{|v_{1}|}+1}\sqrt{\frac{\omega-\Omega}{|v_{1}|}-1}\right]g(\omega)d\omega$ (17)
$Z$ $=i\int_{-\infty}^{\infty}\frac{v_{2}}{|v_{2}|}\left[\frac{\omega-\Omega}{|v_{2}|}\right$.
$\left.\qquad-\sqrt{\frac{\omega-\Omega}{|v_{2}|}+1}\sqrt{\frac{\omega-\Omega}{|v_{2}|}-1}\right]g(\omega)d\omega,$ (18)
where we employ the convention that square roots of a negative real number have a positive imaginary part Abrams et al. (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020). The complex equations (17)–(18) constitute self-consistency relations which allow to determine $r_{1}$, $Z$ and $\Omega$, at least numerically.

### v3

# Finite-size induced random switching of chimeras in a deterministic two-population Kuramoto-Sakaguchi model

@@ Abstract
The two-population Kuramoto-Sakaguchi model for interacting populations of phase oscillators exhibits chimera states whereby one population is synchronised, and the other is desynchronised. Which of the two populations is synchronised depends on the initial conditions. We show that this deterministic model exhibits random switches of their chimera states, alternating between which of the two populations is synchronised and which is not. We show that these random switches are induced by the finite size of the network. We provide numerical evidence that the switches are governed by a Poisson process and that the time between switches grows exponentially with the system size, rendering switches unobservable for all practical purposes in sufficiently large networks. We develop a reduced stochastic model for the synchronised population, based on a central limit theorem controlling the collective effect of the desynchronised population on the synchronised one, and show that this stochastic model well reproduces the statistical behaviour of the full deterministic model. We further determine critical fluctuation sizes capable of inducing switches and provide estimates for the mean switching times from an associated Kramers problem.
@@ I Introduction
Coupled oscillators describe systems ranging from neurons in the brain Crook et al. (1997); Sheeba, Stefanovska, and McClintock (2008); Bhowmik and Shanahan (2012) and synchronous firefly flashing Mirollo and Strogatz (1990), to Josephson junction arrays Watanabe and Strogatz (1994); Wiesenfeld, Colet, and Strogatz (1998) and to power grids Filatrella, Nielsen, and Pedersen (2008); Nishikawa and Motter (2015). The celebrated Kuramoto model and its numerous variants Kuramoto (1984); Sakaguchi and Kuramoto (1986); Strogatz (2000); Pikovsky, Rosenblum, and Kurths (2001); Acebrón et al. (2005); Osipov, Kurths, and Zhou (2007); Arenas et al. (2008); Dörfler and Bullo (2014); Rodrigues et al. (2016) constitute a paradigmatic model to describe the temporal evolution of such coupled oscillators and their often rich dynamical behaviour. Much progress has been made by employing mean-field theory, valid in the thermodynamic limit of infinitely many oscillators Sakaguchi and Kuramoto (1986); Ott and Antonsen (2008); Delgadino, Gvalani, and Pavliotis (2021). However, a number of finite-size effects have recently attracted attention, including finite-size-induced stochastic drift of oscillators that vanishes in the thermodynamic limit Giacomin and Poquet (2015); Luçon (2015); Gottwald (2015, 2017). Such finite-size effects are not amenable to classical mean-field theory and require new analytical approaches Hildebrand, Buice, and Chow (2007); Buice and Chow (2007); Hong, Ha, and Park (2007); Tang (2011); Giacomin and Poquet (2015); Luçon (2015); Gottwald (2015, 2017); OmelchenkoGottwald25 .
@@
We present a dynamical phenomenon that is due entirely to the finite size of the number of oscillators. We describe the emergence of stochastic switching of chimera states in a deterministic system of a two-population Kuramoto-Sakaguchi (KS) model Okuda and Kuramoto (1991); Montbrió, Kurths, and Blasius (2004). Chimeras denote a symmetry-breaking state in which a group of oscillators is synchronised while the rest of the oscillators are desynchronised, and where the partition into a synchronised and desynchronised group depends on the initial condition Kuramoto and Battogtokh (2002); Omel’chenko, Maistrenko, and Tass (2008); Omel’chenko, Wolfrum, and Maistrenko (2010); Omel’chenko (2018). Recently, chimeras were also found in an Ising model with long-range diffusion DeHaroGarciaTorres25 . The existence of chimera states in a two-population KS model is well known and has been studied in either very small networks Panaggio et al. (2016); Burylko, Martens, and Bick (2022) or in the thermodynamic limit Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016). Chimeras can undergo complex dynamical behaviour. In switching (or alternating) chimeras, the synchronised and desynchronised regions alternate in time. To date, such switching has been attributed to either the existence of metastable states Ma, Wang, and Liu (2010); Laing (2012); Buscarino et al. (2015); Semenova et al. (2016); Zhang et al. (2020) or to heteroclinic cycles Bick (2018, 2019); Bick and Lohse (2019); Ebrahimzadeh et al. (2020); Goldschmidt, Pikovsky, and Politi (2019); Brezetsky et al. (2021); Lee and Krischer (2023). In a more complex generalised Kuramoto-Sakaguchi model aperiodic alternating chimeras were observed and analysed that exist in the thermodynamic limit LeeKrischer23b . In contrast, here, we report on effectively random switching of chimeras in a deterministic two-population KS model that is non-transient and mediated entirely by finite-size effects, and that disappears in sufficiently large networks. Numerical simulations will show that the mean switching time increases exponentially with the number of oscillators $N$ and that the variance of the order parameter scales as $1/N$, suggesting an underlying central limit theorem which captures fluctuations around the well-known deterministic thermodynamic limit. We will describe the fluctuations by means of the recently proposed stochastic model reduction for deterministic KS models Yue and Gottwald (2024). We adopt this stochastic modelling framework to parametrise the interaction term that couples the phases of the desynchronised population to those of the synchronised population via a Gaussian process. Gaussian processes are entirely determined by their mean and their covariance function. The mean, represented by the thermodynamic limit of infinitely many oscillators, can be estimated by classical mean field theory. The covariance function will be estimated by first approximating the Gaussian process by a complex Ornstein-Uhlenbeck process, and then numerically determining the parameters of this complex Ornstein-Uhlenbeck process to fit the observed covariance function of the finite-size fluctuations. This modelling strategy captures the statistical features of the full two-population KS model very well, including the distribution of the order parameter and the statistics of the phases of the synchronised oscillators. We will further determine the critical fluctuation size which is sufficient to induce chimera switching, and set up a classical Kramers problem to approximate the mean switching time with reasonable accuracy Kramers (1940). Our modelling strategy requires the system size to be sufficiently large to allow for the underlying central limit theorem to be valid and sufficiently small to prevent the law of large numbers described by the thermodynamic limit and classical mean-field theory.
@@ II Kuramoto-Sakaguchi model for two interacting populations
To account for different populations of oscillators and their mutual interaction, extensions of the KS model were proposed Okuda and Kuramoto (1991); Montbrió, Kurths, and Blasius (2004). The intricate balance between the interactions of oscillators belonging to different populations gives rise to rich dynamical behaviour. In particular, these systems support chimera states in which one population is synchronised while the other is desynchronised Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016); Panaggio et al. (2016); Burylko, Martens, and Bick (2022). Which of the two populations is synchronised and which is not depends on the initial condition. We consider here the dynamics of equally sized populations $P_{1}$ and $P_{2}$ of size $N_{1}=N_{2}=N$, and where each population is governed by the same dynamics in the sense that their native frequencies are drawn from the same distribution and their respective intra- and interpopulation interactions are of equal strength. The dynamics of the phase oscillators are given by
@@
$\frac{d}{dt}\theta_{i}^{(2)}$ $=\omega_{i}^{(2)}+\frac{K}{N}\sum_{j\in P_{2}}\sin(\theta^{(2)}_{j}-\theta^{(2)}_{i}-\lambda)$
@@
Here $\theta_{i}^{(\alpha)}$ denotes the phase of oscillator $i=1,\ldots,N$ in population $\alpha\in\{1,2\}$. The intrapopulation and interpopulation coupling strengths are denoted by $K$ and $\kappa$, respectively. We assume attractive forces and a stronger interaction between oscillators belonging to the same population with $K>\kappa>0$. The parameter $\lambda\in[0,\frac{\pi}{2}]$ accounts for a phase lag. Each oscillator is equipped with a native frequency drawn from a distribution $g(\omega)$. Note that through a change of coordinates into a rotating reference frame, $\theta_{i}^{(\alpha)}(t)\to\theta^{(\alpha)}_{i}(t)-\bar{\omega}\,t$ with mean native frequency $\bar{\omega}=\frac{1}{N}\sum\omega_{i}$, we may assume without loss of generality that the mean native frequency $\bar{\omega}$ is zero. Here we consider normally distributed natural frequencies $g(\omega)\sim\mathcal{N}(0,\sigma_{\omega}^{2})$ for both populations. In all computations, we set $\sigma_{\omega}^{2}=1$, and sample the native frequencies as equiprobable draws from $g(\omega)$. This ensures that both populations have exactly the same set of native frequencies and that no heterogeneity is introduced via random sampling. For independent random draws see Appendix D. We remark that this also best approximates the thermodynamic limit. In the following we shall use greek sub- or superscripts to label populations and roman subscripts to label indices of oscillators.
The two-population KS model (1)–(2) undergoes a transition from desynchronised to synchronised dynamics as the coupling strengths $K$ and $\kappa$ are gradually increased. At lower coupling strengths, the oscillators exhibit desynchronised dynamics and a lack of coherent phase alignment. With increasing coupling strengths, the oscillators align their phases and oscillate with a common rotation frequency $\Omega$. Notably, this synchronisation typically involves both populations achieving coherent phase locking. However, depending on the particular choice of the intrapopulation coupling strength $K$ and the interpopulation coupling strength $\kappa$, the system may exhibit symmetry breaking between the two populations and chimera states appear wherein one population achieves synchronisation while the other remains in a (relatively) desynchronised state Abrams et al. (2008); Laing (2009); Martens, Bick, and Panaggio (2016); Burylko, Martens, and Bick (2022).
The collective behaviour of each of the populations is captured by the complex mean-field variables $z_{\alpha}$, or equivalently by the real mean-field variables $r_{\alpha}$ and $\psi_{\alpha}$ for $\alpha=1,2$, with
$z_{\alpha}\coloneq r_{\alpha}(t)e^{i\psi_{\alpha}(t)}=\frac{1}{N}\sum_{j\in P_{\alpha}}e^{i\theta^{(\alpha)}_{j}(t)}.$ (3)
@@
${\bar{r}}_{\alpha}=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}r_{\alpha}(t)dt,$ (4)
quantifies the degree of synchronisation within a population $P_{\alpha}$; complete phase synchronisation with $\theta^{(\alpha)}_{1}=\theta^{(\alpha)}_{2}=\cdots=\theta^{(\alpha)}_{N}$ implies $\bar{r}_{\alpha}=1$, whereas a fully desynchronised population with $N$ independently and uniformly spread out phases implies $\bar{r}_{\alpha}\sim\mathcal{O}(1/\sqrt{N})$ as a standard Monte Carlo error.
@@ II.1 Mean-field theory
Sakaguchi and Kuramoto Sakaguchi and Kuramoto (1986) developed a mean-field theory for the case of a single population with uniform coupling for the order parameter $r$ and the mean frequency $\Omega$. Mean-field theory has been extended for the two-population KS model (1)–(2) Abrams et al. (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020), which is presented here for completeness. In the two-population KS model, we have mean-field variables $z_{\alpha}$ for each population and the common mean frequency $\Omega$. The two-population KS model (1)–(2) can be written in terms of the complex mean field variables (3) as
$\frac{d}{dt}\theta^{(1)}_{i}$ $=\omega_{i}^{(1)}-\Omega+\textrm{Im}\left[(Kz_{1}+\kappa z_{2})e^{-i\lambda}e^{-i\theta^{(1)}_{i}}\right$], (5)
@@
In view of describing chimera states in which one of the two populations is synchronised, whereas the other one is not, we shall without loss of generality assume that population $P_{1}$ is (partially) synchronised and $P_{2}$ is desynchronised, and we move into the frame of reference of the synchronised cluster with mean phase $\psi_{1}$ and mean frequency $\Omega$. This implies $z_{1}e^{-i\psi_{1}}=r_{1}$ and motivates introducing the phase-shifted order parameter of the desynchronised population
$Z\coloneq z_{2}=r_{2}e^{i\Delta\psi}$ (7)
@@
$\frac{d}{dt}\theta_{j}^{(\alpha)}$ $=\omega_{j}^{(\alpha)}-\Omega+\text{Im}(v_{\alpha}e^{-i\theta_{j}^{(\alpha)}})$ (8)
@@
We focus on stationary chimeras for which, in the thermodynamic limit $\Omega$, $r_{1}$ and $Z$ are constant, corresponding to a stationary distribution of phases. For constant mean-field parameters stationary solutions of (8)–(10) exist, provided the coupling strengths are sufficiently large ensuring
@@
$\theta^{(\alpha)}_{j}$ $=\arcsin\left(\frac{\omega^{(\alpha)}_{j}-\Omega}{|v_{\alpha}|}\right)+\arg(v_{\alpha}).$ (12)
@@
In the thermodynamic limit $N\to\infty$, the phases of oscillators $\theta\in[0,2\pi]$ within population $P_{\alpha}$ with a native frequency of $\omega$ can be described by a probability density function $\rho^{(\alpha)}(\theta,t;\omega)$ for $\alpha=1,2$, satisfying the continuity equation
$\frac{\partial}{\partial t}\rho^{(\alpha)}+\frac{\partial}{\partial\theta}(\nu^{(\alpha)}\rho^{(\alpha)})=0,$ (13)
where the vector fields $\nu^{(\alpha)}(\theta;\omega)$ are given by the right-hand-side of (8) with
$\nu^{(\alpha)}(\theta;\omega)$ $=\omega-\Omega+\text{Im}(v_{\alpha}e^{-i\theta}),$ (14)
where $v_{\alpha}$ is given by (9)–(10). Stationary solutions satisfying $\partial_{t}\rho^{(\alpha)}=0$ depend on the equation parameters. Entrained oscillators with frequencies satisfying (11) are distributed according to the stationary density
@@
$\rho^{(\alpha)}(\theta;\omega)$ $=C^{(\alpha)}(\omega)\frac{1}{\left|\nu^{(\alpha)}(\theta;\omega)\right|}$ (16)
with normalisation constant $C^{(\alpha)}(\omega)=\sqrt{(\omega-\Omega)^{2}-|v^{(\alpha)}|^{2}}/2\pi$, such that the density is high in regions where the phases spend much of their time (cf. (8)–(10)). The order parameters $z_{1}$ and $z_{2}$ are given in the thermodynamic limit as
$z_{1}$ $=\int_{-\infty}^{\infty}\int_{0}^{2\pi}e^{i\theta}\rho^{(1)}(\theta;\omega)g(\omega)d\theta d\omega,$ (17)
$z_{2}$ $=\int_{-\infty}^{\infty}\int_{0}^{2\pi}e^{i\theta}\rho^{(2)}(\theta;\omega)g(\omega)d\theta d\omega.$ (18)
The integration over the native frequencies $\omega$ has to be split into the entrained and non-entrained ranges (cf. (11) for the entrained range), each with their respective stationary density (15) or (16). In particular, upon Fourier decomposition of (16) we obtain
$z_{1}$ $=i\frac{v_{1}}{|v_{1}|}\int_{-\infty}^{\infty}h(\frac{\omega-\Omega}{|v_{1}|})g(\omega)d\omega,$ (19)
$z_{2}$ $=i\frac{v_{2}}{|v_{2}|}\int_{-\infty}^{\infty}h(\frac{\omega-\Omega}{|v_{2}|})g(\omega)d\omega,$ (20)
with
$h(s)$ $=\begin{cases}(1-\sqrt{1-s^{-2}})s&\text{for }|s|>0,\\ s-i\sqrt{1-s^{2}}&\text{for }|s|\leq 1,\end{cases}$ (21)
where we adopt the convenient complex formulation introduced in Omelchenko12 . Note that $v_{\alpha}=v_{\alpha}(z_{1,2})$. Hence, the complex equations (19)–(20) constitute self-consistency relations which allow to determine $z_{1}$, $z_{2}$ and $\Omega$, at least numerically Abrams et al. (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020). We note that the self-consistency relationships (19)–(20) are phase-invariant; therefore, in practice, we additionally impose the restrictions of $z_{1}=r_{1}$ and $z_{2}=Z$ when solving them.